OpenStudy (anonymous):

What is this crap? I don't understand. Create your own equation written in standard form. Determine any point that is a solution and will be on this line. Justify your answer algebraically.

OpenStudy (anonymous):

@Luigi0210 HALP :d

OpenStudy (luigi0210):

Linear equation? or polynomial? etc.?

OpenStudy (anonymous):

Linear equation thats what lesson thing its in

OpenStudy (anonymous):

Is this for Algebra 2 ?

OpenStudy (anonymous):

Alg 1

OpenStudy (anonymous):

Oh okay . Do you know the formula ?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

@kdog771998 can u help

OpenStudy (luigi0210):

If it's a linear, then standard form is something like: Ax+By=C And @kdog771998 that's slope-intercept form

OpenStudy (anonymous):

Lol oppsss .

OpenStudy (anonymous):

i have no clue so i just want anyone to help :d

OpenStudy (anonymous):

I was confused with algebra 2 standard form thats ax^2 + bx + c

OpenStudy (luigi0210):

For quadratic equations, yes.

OpenStudy (anonymous):

Okay so then the formulas Ax + By=C

OpenStudy (anonymous):

so could the equation be 10x + 14b = c?

OpenStudy (shamil98):

No.

OpenStudy (luigi0210):

C is just a constant, any real number.

OpenStudy (shamil98):

Ax+By = C A and B are coefficients, x and y are variables, C is a constant.

OpenStudy (luigi0210):

Dang it sham, how did you get here? ლ(■_■ლ)

OpenStudy (shamil98):

In standard form you want all your values in whole numbers, (so no decimals/fractions).

OpenStudy (shamil98):

For example. 4x + 5y = 10 This is an example of line in standard-form.

OpenStudy (shamil98):

if you had something like 2/3 x + 2/3 y = 2 you would want to multiply the equation by 3 to get whole numbers 3(2/3x+2/3y=2) becomes 2x + 2y = 6

OpenStudy (anonymous):

Ok so we have the equation right?

OpenStudy (anonymous):

so what would i need to do next to get my answer?

OpenStudy (anonymous):

for the second half

OpenStudy (shamil98):

Let's use 2x + 2y = 6 What values make this equation true? Let's make y = 0 so 2x + 2(0) = 6 2x = 6 x = 3 so one solution is (3,0)

OpenStudy (shamil98):

Another solution could be when x = 0 so 2(0) + 2y = 6 2y = 6 y = 3 so (0,3) is another solution

OpenStudy (anonymous):

Ohh i get it now, so you sub both for 0 to get the points. Thanks sham